# TPU as Cryptographic Accelerator

### v2

@@ Abstract
In the last decades, a large number of cryptographic tools with fascinating features have been developed, such as fully homomorphic encryption (FHE) [4], zero-knowledge proof (ZKP) [8], and post-quantum cryptography schemes (PQC) [3]. FHE allows a third party to conduct arbitrary computation (addition and multiplication) on ciphertexts and ZKP helps a prover to convince a verifier on the correctness of a statement without disclosing the secret involved in the statement. These two types of cryptographic schemes can be utilized to build many interesting security applications, such as secure AI inferencing [7, 15] and privacy-preserving cryptocurrency exchange [31]. One major obstacle that hinders the wide adoption of these schemes is their high computation cost. PQC schemes are designed to resist attacks with quantum computers. They are not as expensive as FHE and ZKP, but all fall into the category of asymmetric cryptography (digital signature and public key encryption/key encapsulation mechanism) and are computation-intensive by nature.
@@
Many works have been done on improving the polynomial multiplication from the algorithmic perspective, such as Karatsuba multiplication [16] and number theoretic transform (NTT) based multiplication [20]. Implementation, especially hardware-based polynomial multiplication implementation, is another important research direction. The two major hardware platforms that are studied are GPU and Field Programmable Gate Array (FPGA). Both GPU and FPGA support parallel computing, which is utilized to accelerate polynomial multiplication [30, 21].
One emerging computation hardware that is largely ignored is the tensor processing unit (TPU), or neural processing unit (NPU). These two names usually refer to the same type of hardware, and we use the term TPU in the rest of the paper. The concept of TPU was originally proposed by Google to meet the explosive computation power demand of artificial intelligence (AI) applications and has been proven to be a big success [24]. Many TPU designs have been proposed since then, especially for the edge computing scenario. Compared with GPU and FPGA, TPU has its unique advantages. On the one hand, it is closer to Application-Specific Integrated Circuits (ASIC) and can significantly accelerate specific operations with low energy costs. On the other hand, TPU still provides a certain level of flexibility and allows end users to program it.
While TPUs are widely used in AI [10], their application in cryptography is largely unexplored. To the best of our knowledge, this is the first work investigating TPUs for polynomial multiplication, a core operation in FHE, ZKP, and PQC. To leverage TPUs’ MACs, we convert polynomial multiplications to matrix multiplications. However, TPU’s AI-centric design struggles with these matrices. We address this using a carefully chosen RNS [6] and matrix partitioning. Our prototype on Google’s cloud TPU demonstrates TPU’s promise for cryptographic acceleration, with significant room for improvement given the limitations of current TPU software tools.
@@
Polynomials are used intensively in various public key cryptographic schemes mainly for two reasons: (i) Polynomial is related to popular hard problems used in cryptography. Many PQC and FHE schemes are built atop ring LWE problems [29], and using a polynomial ring is a major way to instantiate a ring LWE instance. (ii) Polynomial is an effective way for computation representation. An effective computation representation is critical for ZKP, and many ZKP schemes use polynomials in one way or another to represent the computation (or the statement) the prover needs to prove. Polynomial operations, especially polynomial multiplications, usually comprise a large portion of the computation cost of these cryptographic schemes and are the performance bottleneck.
@@
The dedicated and optimized circuit is usually much faster and more power-efficient than CPU and GPU on MAC operations. According to [13], a Haswell server equipped with the original version of Google TPU can achieve 92 TOPS/s with a TDP of 861W, while the server itself can only achieve 2.6 TOPS/s with a TDP of 504W. After the first generation, Google has released three revised versions of TPU and the latest version is V4 [12]. A newer version usually has more Matrix Multiply Units and memory on a single device.
@@ 2.1 Polynomials Used in Cryptography and Their Operations
where $a_{i}\in\mathbb{Z}_{q},i=0,\ldots,n-1$. The two operations defined on the ring are polynomial addition and polynomial multiplication. Polynomial addition is straightforward, i.e., two corresponding coefficients of two polynomials are added together and then reduced modulo $q$. Multiplication is more complex and consumes most of the computation time. We review two common ways to do multiplication, which are closely related to the representation of polynomials: (i) Multiplication with coefficient representation.Coefficient representation is a natural way to store a polynomial, as demonstrated in Eq. (1). When two polynomials defined in $\mathbb{Z}_{q}[x]/(x^{n}+1)$ are stored using coefficient representation, a straightforward approach is first multiplying them as if they are defined in $\mathbb{Z}[x]$, and then reducing the polynomial by taking modulo $x^{n}+1$ and $q$ to obtain the final result. One disadvantage of this approach is that the intermediate result has a degree of at most $22n-$ 2 with coefficients as large as $(q-1)^{2}$. We may integrate these modulo operations into the multiplication to save some time. The complexity of a straightforward multiplication is $O(n^{2})$, and using some techniques like Karatsuba algorithm can bring down the complexity to the level of $O(n^{\log_{2}3})$ [16]. (ii) Multiplication with value representation.Another common way to store a polynomial is using value representation. For the polynomial $f(x)$ given in Eq. (1) represented with coefficients, one can evaluate it at $n$ different values for $x$, and store all these $n$ pairs $[x_{1},y_{1}=f(x_{1})],[x_{2},y_{2}=f(x_{2})],\ldots,[x_{n},y_{n}=f(x_{n})]$. If we always use $x_{i}=i\in\mathbb{Z},i=1,2,\ldots,n$ for evaluating a polynomial, $x_{i}$ can be ignored and only $y_{i}$s need to be stored (we assume $n<q$, which usually holds). These pairs (or $y_{i}$s) fully determine the polynomial and we can convert them back to coefficient representation using Lagrange interpolation. The multiplication of polynomials stored using value representation is easy, i.e., we only need to perform an element-wise multiplication. As most cryptography schemes use polynomials with coefficient representation, one needs to convert them to value representations first. A straightforward way of conversion is evaluating a polynomial at each given value, and the complexity of each evaluation is $O(n)$. Since the polynomial is evaluated at $n$ values, the overall complexity is $O(n^{2})$, which is more expensive than the multiplication operation in value representation. Because the multiplication result may need to be involved in other cryptographic operations, we usually need to convert it back to coefficient representation. This inverse operation can be done using straightforward Lagrange interpolation, and its complexity is also $O(n^{2})$. When the two parameters $q$ and $n$ of $\mathbb{Z}_{q}[x]/(x^{n}+1)$ meet certain requirements, the conversion and reverse conversion can leverage number theoretic transform (NTT), and the complexity is reduced to the level of $O(n\log n)$.
@@
Because of the lower computation complexity, multiplication using value representation receives a lot of attention. However, this approach has several limitations: (i) Size of the problem.The lower computation complexity does not guarantee faster execution in practice. NTT is faster only for relatively large polynomial degrees. When the cryptography schemes use polynomials with low degrees, the difference between these two multiplication methods is small, and value representation can be even slower. (ii) Restrictions on parameters of $\mathbb{Z}_{q}[x]/(x^{n}+1)$.The low computation complexity of multiplication using value representation relies on the applicability of NTT and requires the polynomials to be NTT-friendly (e.g., $n$ is a power of $2$ and $q\equiv 1\mod n$). While there are some works on losing these requirements [5], it is hard to get rid of them completely. These requirements restrict the parameter selection for the cryptography scheme, which is not desirable. (iii) Compatibility with the hardware processor.Value representation-based multiplication can be implemented as software but it does not guarantee that the software can fully utilize the underlying hardware capability.
@@ 2.3 Challenges of Using TPU for Polynomial Multiplication
The way of converting polynomial multiplication to matrix operation described above is not enough for utilizing TPU for cryptographic operations. The first challenge is the degree of the polynomials. For polynomials with degree $n$, the corresponding matrix dimension is $n\times n$. The degree of the polynomials involved in a typical cryptographic scheme can be large, and the corresponding matrix cannot fit the TPU hardware for computation. The second challenge is the size of the polynomial coefficients. Most TPUs are designed for machine learning tasks, and a significant amount of work has been done to reduce the size of values involved in a machine learning model. For instance, the Google TPU supports bf16 natively, which is a 16-bit floating point number system and has lower precision compared with the standard 32-bit floating point number, but provides similar model performance [27]. However, polynomials used in cryptography usually have relatively large coefficients, from tens of bits to a few hundred bits. Furthermore, these coefficients are generally defined on a finite and discrete algebra structure, which is not natively consistent with the data type supported by the TPU. Table 1 summarizes some polynomial parameters used by different cryptographic schemes. We describe the way to overcome these two challenges in the following sections.
@@ 2.4 Handling Large Coefficients
Classical digital computers store and process values presented using a limited number of bits. For large values, there are two strategies to process them: (i) Using an approximated value to replace the original one.With this approach, limited storage is used to store a value that is close to the real value. Floating point number schemes like IEEE 754 [14] and bfloat16 [27] fall into this category of solution. This strategy is effective for a wide range of applications where the problems do not require 100% computation accuracy. (ii) Using multiple basic units to represent the original one.The other one is breaking a large value into several parts so each part can be stored and processed by the hardware. Accordingly, an operation on the original large values is also decomposed into a sequence of operations on the small units, which can be handled by the hardware directly. One common example is big integer operations, where an integer with more than 32 or 64 bits is stored in multiple basic data structures, and dedicated algorithms are implemented to support the common arithmetic operations [2]. This strategy is widely used in cryptographic libraries to support schemes such as RSA.
@@
To overcome this challenge, we propose to utilize the residue number system (RNS) [26] for coefficients representation and computation. A residue number system is defined by a set of $k$ co-prime integers called the moduli $\{m_{1},m_{2},\ldots,m_{k}\}$. An integer $x$ is represented in the RNS by its remainder $[x_{1},x_{2},\ldots,x_{k}]$, where $x_{i}\equiv x\mod m_{i}$ for every $i$.
@@ 2.7 Putting All Pieces Together
Because of the popularity of AI applications, there is a huge interest in TPU hardware design and development [25]. However, there are only a limited number of TPU products that are available for general developers. One of the most popular ones is Google TPU, which can be conveniently accessed through Google Cloud. In this work, we mainly evaluate the design using Google cloud-based TPU.
Overview of the Google cloud-based TPU. Google provides two avenues to access its TPU hardware: creating a TPU-equipped virtual machine (VM) in the cloud, or utilizing Colab, an online development environment with TPU connectivity. Both options support TensorFlow, PyTorch, and JAX (https://github.com/google/jax) for TPU software development, all Python-based and originally tailored for AI applications.
For our prototyping and benchmarking, we opted for the TPU VM in conjunction with JAX. This combination provides the most flexibility, allowing us to configure the TPU VM with different hardware versions and topologies. JAX also facilitates non-AI computations and offers directives for program optimization. Currently, Google Colab only permits requesting a TPU instance without further hardware configuration. Moreover, Colab lacks support for JAX 0.4, which we utilize in our experiments.
@@
Table 2: Description of the benchmarking environments in the Google cloud.
Environment Instance Description
Env 1 (Figure 2(a)) Ubuntu 22.04, TPU v2-8, which includes eight TPU v2 physical cores and 64 GB memory.
Env 2 (Figure 2(b)) Ubuntu 22.04, TPU v3-8, which includes eight TPU v3 physical cores and 128 GB memory.
Figure 2(a) and Figure 2(b) summarize the benchmarking results using two execution environments Env 1 and Env 2 respectively. For each figure, we fix the value of $m$ and the x-axis is different values of $n$. The y-axis is the execution time and the unit is the second. We observe several things from the experiment results: (i) For small matrix dimensions, the computation time is a constant value. After the threshold, the computation time increases linearly with the value of $n$. Note that the overall computation complexity is $O(m\times n\times n)$ (if the computation is done using the straightforward algorithm), which is not consistent with the experiment results. One potential reason is that the TPU architecture carries out the computation in a highly parallel manner and reduces the latency. (ii) The slopes of curves for different $m$ values are similar, which means that $m$ does not affect the cost much. The potential reason is that the TPU hardware supports $128\times 128$ matrix multiplication, and the largest $m$ we choose is within this range. (iii) TPU v3 is faster than TPU v2, but the difference is not significant, especially for cases with small parameters. This may be because both versions have the same number of matrix multiplication hardware (as described in Table 2).
Limitations of the current benchmarking. From the experiment results, we can see that TPU is a promising cryptography accelerator. The major limitation of current benchmarking is that we cannot control the execution in detail and it is possible the potential of TPUs is not fully utilized. For instance, TPU v3 has a much larger memory but the current programming environment does not allow us to control it directly. When multiple matrix multiplications are required, it is possible to create a pipeline to reduce the overall latency. But Google TPU does not support such fine-tuning at this time.
@@
Karatsuba multiplication. The idea of Karatsuba multiplication [16] is trading one expensive multiplication with multiple cheap addition/subtraction operations. For two polynomials $f(x)$ and $g(x)$ with degree $n$, we can rewrite them as $f(x)=f_{1}(x)x^{m}+f_{0}(x)$ and $g(x)=g_{1}(x)x^{m}+g_{0}(x)$, where $f_{1},g_{1}$ are polynomials of degree $n-m$, and $f_{0},g_{0}$ are polynomials of degree $m-1$. The product is then
@@
This technique can be adopted to accelerate TPU-based polynomial multiplication too. Specifically, the original polynomials are decomposed into multiple ones with a lower degree, and then the TPU is utilized to calculate multiplications. This method can be leveraged to replace the high-degree polynomial handling approach described in Section 2.5. A key challenge is TPU’s inefficiency for polynomial addition/subtraction. Employing Karatsuba multiplication might necessitate additional hardware, incurring communication overhead and complicating optimization.
Fine tuning the design. In this work, we mainly address the feasibility issue of using TPU for polynomial multiplication. In practice, a cryptographic scheme usually compromises multiple polynomial multiplications in a specific order. There are several strategies that can be considered to fine-tune the computation: (i) Fixing one operand polynomial as much as possible.One operand polynomial needs to be converted to the matrix form to incorporate the modulo operation into multiplication. The creation of the matrix and the process of loading the matrix into the TPU memory are expensive compared with the matrix multiplication computation. Therefore, it is better to reuse a generated and loaded matrix to minimize the preparation cost. (ii) Optimize the computation and I/O.Polynomial multiplication in cryptography is both computation and I/O intensive, and the working environment is complex. As TPU works as a passive accelerator, data needs to be transferred from the host machine to a TPU card. As a machine usually has multiple TPUs, we also need to consider which TPU should be used at a given time. It is possible that the intermediate result produced by the TPU needs to be transferred back to the host to process and we want to make sure the TPU is busy with other tasks during this process.
@@
Accelerating multiplication of polynomials defined in specific rings has been studied from different perspectives. We summarize these works and discuss their relationships with our work in this section.
Algorithmic approaches. Multiplication acceleration methods are studied for both coefficient and value representations, e.g., Karatsuba algorithm [16] and various NTT approaches [5]. Some works also consider more efficient modulo operation, which is heavily involved in polynomial operations [28]. These works are orthogonal to our work and have the potential to be utilized to further improve the polynomial multiplication performance with the TPU.
Hardware-oriented optimization approach. This line of research investigates the utilization of existing hardware features to accelerate polynomial multiplication on specific rings. Several types of hardware are studied for this purpose, including CPU [1], GPU [23], and FPGA [22]. Compared with other types of hardware, TPU provides a new balance between flexibility, power consumption, and parallelism. Table 3 compares TPU with other computation platforms from the perspectives of flexibility, power consumption, and parallelism.
@@
TPU related works. Most existing works on TPU focus on its architecture [13, 11] and applications in AI [18, 17], with a few exceptions considering TPU for non-AI use cases. For instance, Hsu and Tseng developed a tool to use edge TPU for general-purpose computation [9], and Lewis et.al discussed the use of TPU for large-scale linear algebra computation [19]. None of these works consider the special requirements of cryptographic schemes.
@@
Accelerating polynomial multiplication defined in specific algebra structures is crucial for improving the performance of a wide range of cryptographic schemes. In this work, we investigate the application of the emerging AI processor (TPU) for polynomial multiplication to accelerate corresponding cryptographic operations. We address two major challenges in this process, i.e., using RNS to handle large coefficients that are not supported by the TPU and adopting a divide-and-conquer strategy to process polynomials with a high degree that cannot fit into hardware. We conduct preliminary experiments to verify the feasibility of the proposed solution and discuss possible ways to further improve the performance.
For future works, we plan to work on three items: (i) Investigating the method of applying TPU for NTT-based polynomial multiplication; (ii) Further improve the prototype by optimizing the hardware configuration for specific polynomial rings; and (iii) Design and verify end-to-end cryptographic schemes with TPU and investigate corresponding optimization methods.

### v3

@@ Abstract
In the last decades, a large number of cryptographic tools with fascinating features have been developed, such as fully homomorphic encryption (FHE) [Brakerski(2012)], zero-knowledge proof (ZKP) [Groth(2016)], and post-quantum cryptography schemes (PQC) [Bos et al.(2018)Bos, Ducas, Kiltz, Lepoint, Lyubashevsky, Schanck, Schwabe, Seiler, and Stehlé]. FHE allows a third party to conduct arbitrary computation (addition and multiplication) on ciphertexts and ZKP helps a prover to convince a verifier on the correctness of a statement without disclosing the secret involved in the statement. These two types of cryptographic schemes can be utilized to build many interesting security applications, such as secure AI inferencing [Gilad-Bachrach et al.(2016)Gilad-Bachrach, Dowlin, Laine, Lauter, Naehrig, and Wernsing, Kang et al.(2022)Kang, Hashimoto, Stoica, and Sun] and privacy-preserving cryptocurrency exchange [Xu et al.(2020)Xu, Chen, Gao, Kasichainula, Fernandez, Carbunar, and Shi]. One major obstacle that hinders the wide adoption of these schemes is their high computation cost. PQC schemes are designed to resist attacks with quantum computers. They are not as expensive as FHE and ZKP, but all fall into the category of asymmetric cryptography (digital signature and public key encryption/key encapsulation mechanism) and are computation-intensive by nature.
@@
Many works have been done on improving the polynomial multiplication from the algorithmic perspective, such as Karatsuba multiplication [Karatsuba and Ofman(1962)] and number theoretic transform (NTT) based multiplication [Liang and Zhao(2022)]. Implementation, especially hardware-based polynomial multiplication implementation, is another important research direction. The two major hardware platforms that are studied are GPU and Field Programmable Gate Array (FPGA). Both GPU and FPGA support parallel computing, which is utilized to accelerate polynomial multiplication [Shivdikar et al.(2022)Shivdikar, Jonatan, Mora, Livesay, Agrawal, Joshi, Abellán, Kim, and Kaeli, Mert et al.(2019)Mert, Öztürk, and Savaş].
One emerging computation hardware that is largely ignored is the tensor processing unit (TPU), or neural processing unit (NPU). These two names usually refer to the same type of hardware, and we use the term TPU in the rest of the paper. The concept of TPU was originally proposed by Google to meet the explosive computation power demand of artificial intelligence (AI) applications and has been proven to be a big success [Norrie et al.(2021)Norrie, Patil, Yoon, Kurian, Li, Laudon, Young, Jouppi, and Patterson]. Many TPU designs have been proposed since then, especially for the edge computing scenario. Compared with GPU and FPGA, TPU has its unique advantages. On the one hand, it is closer to Application-Specific Integrated Circuits (ASIC) and can significantly accelerate specific operations with low energy costs. On the other hand, TPU still provides a certain level of flexibility and allows end users to program it.
While TPUs are widely used in AI [Huot et al.(2019)Huot, Chen, Clapp, Boneti, and Anderson], their application in cryptography is largely unexplored. To the best of our knowledge, this is the first work investigating TPUs for polynomial multiplication, a core operation in FHE, ZKP, and PQC. To leverage TPUs’ MACs, we convert polynomial multiplications to matrix multiplications. However, TPU’s AI-centric design struggles with these matrices. We address this using a carefully chosen RNS [Garner(1959)] and matrix partitioning. Our prototype on Google’s cloud TPU demonstrates TPU’s promise for cryptographic acceleration, with significant room for improvement given the limitations of current TPU software tools.
@@
Polynomials are used intensively in various public key cryptographic schemes mainly for two reasons: (i) Polynomial is related to popular hard problems used in cryptography. Many PQC and FHE schemes are built atop ring LWE problems [Rosca et al.(2018)Rosca, Stehlé, and Wallet], and using a polynomial ring is a major way to instantiate a ring LWE instance. (ii) Polynomial is an effective way for computation representation. An effective computation representation is critical for ZKP, and many ZKP schemes use polynomials in one way or another to represent the computation (or the statement) the prover needs to prove. Polynomial operations, especially polynomial multiplications, usually comprise a large portion of the computation cost of these cryptographic schemes and are the performance bottleneck.
@@
The dedicated and optimized circuit is usually much faster and more power-efficient than CPU and GPU on MAC operations. According to [Jouppi et al.(2018)Jouppi, Young, Patil, and Patterson], a Haswell server equipped with the original version of Google TPU can achieve 92 TOPS/s with a TDP of 861W, while the server itself can only achieve 2.6 TOPS/s with a TDP of 504W. After the first generation, Google has released three revised versions of TPU and the latest version is V4 [Jouppi et al.(2021)Jouppi, Yoon, Ashcraft, Gottscho, Jablin, Kurian, Laudon, Li, Ma, Ma, et al.].
@@ 2.1 Polynomials Used in Cryptography and Their Operations
where $a_{i}\in\mathbb{Z}_{q},i=0,\ldots,n-1$. The two operations defined on the ring are polynomial addition and polynomial multiplication. Polynomial addition is straightforward, i.e., two corresponding coefficients of two polynomials are added together and then reduced modulo $q$. Multiplication is more complex and consumes most of the computation time. We review two common ways to do multiplication, which are closely related to the representation of polynomials: (i) Multiplication with coefficient representation.Coefficient representation is a natural way to store a polynomial, as demonstrated in Eq. (1). When two polynomials defined in $\mathbb{Z}_{q}[x]/(x^{n}+1)$ are stored using coefficient representation, a straightforward approach is first multiplying them as if they are defined in $\mathbb{Z}[x]$, and then reducing the polynomial by taking modulo $x^{n}+1$ and $q$ to obtain the final result. One disadvantage of this approach is that the intermediate result has a degree of at most $22n-$ 2 with coefficients as large as $(q-1)^{2}$. We may integrate these modulo operations into the multiplication to save some time. The complexity of a straightforward multiplication is $O(n^{2})$, and using some techniques like Karatsuba algorithm can bring down the complexity to the level of $O(n^{\log_{2}3})$ [Karatsuba and Ofman(1962)]. (ii) Multiplication with value representation.Another common way to store a polynomial is using value representation. For the polynomial $f(x)$ given in Eq. (1) represented with coefficients, one can evaluate it at $n$ different values for $x$, and store all these $n$ pairs $[x_{1},y_{1}=f(x_{1})],[x_{2},y_{2}=f(x_{2})],\ldots,[x_{n},y_{n}=f(x_{n})]$. If we always use $x_{i}=i\in\mathbb{Z},i=1,2,\ldots,n$ for evaluating a polynomial, $x_{i}$ can be ignored and only $y_{i}$s need to be stored (we assume $n<q$, which usually holds). These pairs (or $y_{i}$s) fully determine the polynomial and we can convert them back to coefficient representation using Lagrange interpolation. The multiplication of polynomials stored using value representation is easy, i.e., we only need to perform an element-wise multiplication. As most cryptography schemes use polynomials with coefficient representation, one needs to convert them to value representations first. A straightforward way of conversion is evaluating a polynomial at each given value, and the complexity of each evaluation is $O(n)$. Since the polynomial is evaluated at $n$ values, the overall complexity is $O(n^{2})$, which is more expensive than the multiplication operation in value representation. Because the multiplication result may need to be involved in other cryptographic operations, we usually need to convert it back to coefficient representation. This inverse operation can be done using straightforward Lagrange interpolation, and its complexity is also $O(n^{2})$. When the two parameters $q$ and $n$ of $\mathbb{Z}_{q}[x]/(x^{n}+1)$ meet certain requirements, the conversion and reverse conversion can leverage number theoretic transform (NTT), and the complexity is reduced to the level of $O(n\log n)$.
@@
Because of the lower computation complexity, multiplication using value representation receives a lot of attention. However, this approach has several limitations: (i) Size of the problem.The lower computation complexity does not guarantee faster execution in practice. NTT is faster only for relatively large polynomial degrees. When the cryptography schemes use polynomials with low degrees, the difference between these two multiplication methods is small, and value representation can be even slower. (ii) Restrictions on parameters of $\mathbb{Z}_{q}[x]/(x^{n}+1)$.The low computation complexity of multiplication using value representation relies on the applicability of NTT and requires the polynomials to be NTT-friendly (e.g., $n$ is a power of $2$ and $q\equiv 1\mod n$). While there are some works on losing these requirements [Chung et al.(2021)Chung, Hwang, Kannwischer, Seiler, Shih, and Yang], it is hard to get rid of them completely. These requirements restrict the parameter selection for the cryptography scheme, which is not desirable. (iii) Compatibility with the hardware processor.Value representation-based multiplication can be implemented as software but it does not guarantee that the software can fully utilize the underlying hardware capability.
@@ 2.3 Challenges of Using TPU for Polynomial Multiplication
The way of converting polynomial multiplication to matrix operation described above is not enough for utilizing TPU for cryptographic operations. The first challenge is the degree of the polynomials. For polynomials with degree $n$, the corresponding matrix dimension is $n\times n$. The degree of the polynomials involved in a typical cryptographic scheme can be large, and the corresponding matrix cannot fit the TPU hardware for computation. The second challenge is the size of the polynomial coefficients. Most TPUs are designed for machine learning tasks, and a significant amount of work has been done to reduce the size of values involved in a machine learning model. For instance, the Google TPU supports bf16 natively, which is a 16-bit floating point number system and has lower precision compared with the standard 32-bit floating point number, but provides similar model performance [Osorio et al.(2022)Osorio, Armejach, Petit, Henry, and Casas]. However, polynomials used in cryptography usually have relatively large coefficients, from tens of bits to a few hundred bits. Furthermore, these coefficients are generally defined on a finite and discrete algebra structure, which is not natively consistent with the data type supported by the TPU. Table 1 summarizes some polynomial parameters used by different cryptographic schemes. We describe the way to overcome these two challenges in the following sections.
@@ 2.4 Handling Large Coefficients
Classical digital computers store and process values presented using a limited number of bits. For large values, there are two strategies to process them: (i) Using an approximated value to replace the original one.With this approach, limited storage is used to store a value that is close to the real value. Floating point number schemes like IEEE 754 [Kahan(1996)] and bfloat16 [Osorio et al.(2022)Osorio, Armejach, Petit, Henry, and Casas] fall into this category of solution. This strategy is effective for a wide range of applications where the problems do not require 100% computation accuracy. (ii) Using multiple basic units to represent the original one.The other one is breaking a large value into several parts so each part can be stored and processed by the hardware. Accordingly, an operation on the original large values is also decomposed into a sequence of operations on the small units, which can be handled by the hardware directly. One common example is big integer operations, where an integer with more than 32 or 64 bits is stored in multiple basic data structures, and dedicated algorithms are implemented to support the common arithmetic operations [Biswas et al.(2003)Biswas, Bandyopadhyay, and Banerjee]. This strategy is widely used in cryptographic libraries to support schemes such as RSA.
@@
To overcome this challenge, we propose to utilize the residue number system (RNS) [Omondi and Premkumar(2007)] for coefficients representation and computation. A residue number system is defined by a set of $k$ co-prime integers called the moduli $\{m_{1},m_{2},\ldots,m_{k}\}$. An integer $x$ is represented in the RNS by its remainder $[x_{1},x_{2},\ldots,x_{k}]$, where $x_{i}\equiv x\mod m_{i}$ for every $i$.
@@ 2.7 Putting All Pieces Together
Due to the rising popularity of AI applications [Karanjai et al.(2024)Karanjai, Xu, and Shi], there is a growing interest in the design and development of TPU hardware [Olsen(2017)]. However, the number of TPU products accessible to general developers remains limited. One of the most widely used options is Google’s TPU, which is what we primarily evaluate our design using.
Overview of Google Cloud-based TPU. Google offers two primary methods for accessing its TPU hardware: creating a TPU-equipped virtual machine (VM) in the cloud or utilizing Colab, an online development platform with TPU connectivity. Both approaches support popular frameworks such as TensorFlow, PyTorch, and JAX (https://github.com/google/jax), which are all Python-based and originally designed for AI applications.
For our prototyping and benchmarking, we chose the TPU VM paired with JAX, as it provides the greatest flexibility. This setup allows us to configure the TPU VM with different hardware versions and topologies. JAX also supports non-AI computations and includes tools for optimizing program performance. Currently, Google Colab only allows requesting a TPU instance without additional hardware configuration options. Furthermore, Colab does not yet support JAX 0.4, which we used in our experiments.
@@
Table 2: Benchmarking environments on Google Cloud.
Env Description (Physical Cores)
Env 1 (Figure 2(a)) Ubuntu 22.04, TPU v2-8: 8 cores, 64 GB memory.
Env 2 (Figure 2(b)) Ubuntu 22.04, TPU v3-8: 8 cores, 128 GB memory.
Figure 2(a) and Figure 2(b) summarize the benchmarking results using two execution environments Env 1 and Env 2 respectively. For each figure, we fix the value of $m$ and the x-axis is different values of $n$. The y-axis is the execution time and the unit is the second. We observe several things from the experiment results: (i) For small matrix dimensions, the computation time is a constant value. After the threshold, the computation time increases linearly with the value of $n$. Note that the overall computation complexity is $O(m\times n\times n)$ (if the computation is done using the straightforward algorithm), which is not consistent with the experiment results. One potential reason is that the TPU architecture carries out the computation in a highly parallel manner and reduces the latency. (ii) The slopes of curves for different $m$ values are similar, which means that $m$ does not affect the cost much. The potential reason is that the TPU hardware supports $128\times 128$ matrix multiplication, and the largest $m$ we choose is within this range. (iii) TPU v3 is faster than TPU v2, but the difference is not significant, especially for cases with small parameters. This may be because both versions have the same number of matrix multiplication hardware (Table 2).
Limitations of the current benchmarking.
The experimental results demonstrate that TPUs hold significant promise as cryptography accelerators. However, the primary limitation of our current benchmarking lies in the inability to control execution at a granular level, potentially leaving TPU capabilities underutilized. For instance, while TPU v3 offers substantially larger memory, the current programming environment does not provide direct control over its usage. When multiple matrix multiplications are required, creating a pipeline could reduce overall latency. Unfortunately, Google’s TPU framework does not yet support such fine-tuning.
@@
Karatsuba multiplication. The idea of Karatsuba multiplication [Karatsuba and Ofman(1962)] is trading one expensive multiplication with multiple cheap addition/subtraction operations. For two polynomials $f(x)$ and $g(x)$ with degree $n$, we can rewrite them as $f(x)=f_{1}(x)x^{m}+f_{0}(x)$ and $g(x)=g_{1}(x)x^{m}+g_{0}(x)$, where $f_{1},g_{1}$ are polynomials of degree $n-m$, and $f_{0},g_{0}$ are polynomials of degree $m-1$. The product is then
@@
This technique can also be applied to accelerate TPU-based polynomial multiplication. Specifically, the original polynomials are decomposed into multiple lower-degree polynomials, and the TPU is then employed to perform the multiplications. This method can be used in place of the high-degree polynomial handling approach described in Section 2.5. However, a key challenge lies in the TPU’s inefficiency for polynomial addition and subtraction. The use of Karatsuba multiplication may require additional hardware, introducing communication overhead and complicating optimization.
Fine-tuning the design. In this work, we primarily address the feasibility of using TPUs for polynomial multiplication. In practice, cryptographic schemes often require multiple polynomial multiplications in a specific sequence. Several strategies can be considered to fine-tune the computation: (i) Fixing one operand polynomial as much as possible.One operand polynomial needs to be converted to matrix form to incorporate the modulo operation into the multiplication. Creating the matrix and loading it into TPU memory is costly compared to matrix multiplication itself. Therefore, reusing a pre-generated and loaded matrix minimizes the preparation overhead. (ii) Optimizing computation and I/O.Polynomial multiplication in cryptography is both computation- and I/O-intensive, and the environment is complex. As TPUs function as passive accelerators, data must be transferred from the host machine to the TPU card. Since a machine typically has multiple TPUs, selecting the appropriate TPU at a given time is crucial. Moreover, intermediate results generated by the TPU might need to be sent back to the host for further processing, so ensuring the TPU remains occupied with other tasks during this transfer is essential.
@@
Accelerating multiplication of polynomials defined in specific rings has been studied from different perspectives.
Algorithmic approaches. Multiplication acceleration methods are studied for both coefficient and value representations, e.g., Karatsuba algorithm [Karatsuba and Ofman(1962)] and various NTT approaches [Chung et al.(2021)Chung, Hwang, Kannwischer, Seiler, Shih, and Yang]. Some works also consider more efficient modulo operation, which is heavily involved in polynomial operations [Posch and Posch(1995)]. These works are orthogonal to our work and have the potential to be utilized to further improve the polynomial multiplication performance with the TPU.
Hardware-oriented optimization approach. This line of research investigates the utilization of existing hardware features to accelerate polynomial multiplication on specific rings. Several types of hardware are studied for this purpose, including CPU [Aguilar-Melchor et al.(2016)Aguilar-Melchor, Barrier, Guelton, Guinet, Killijian, and Lepoint], GPU [Ni and Zhu(2023)], and FPGA [Mert et al.(2020)Mert, Öztürk, and Savaş]. Compared with other types of hardware, TPU provides a new balance between flexibility, power consumption, and parallelism. Table 3 compares TPU with other computation platforms.
@@
TPU related works. Most existing works on TPU focus on its architecture [Jouppi et al.(2018)Jouppi, Young, Patil, and Patterson, Jouppi et al.(2023)Jouppi, Kurian, Li, Ma, Nagarajan, Nai, Patil, Subramanian, Swing, Towles, et al.] and applications in AI [Kumar et al.(2019)Kumar, Bitorff, Chen, Chou, Hechtman, Lee, Kumar, Mattson, Wang, Wang, et al., Kimm et al.(2021)Kimm, Paik, and Kimm], with a few exceptions considering TPU for non-AI use cases. For instance, Hsu and Tseng developed a tool to use edge TPU for general-purpose computation [Hsu and Tseng(2021)], and Lewis et.al discussed the use of TPU for large-scale linear algebra computation [Lewis et al.(2022)Lewis, Beall, Ganahl, Hauru, Mallick, and Vidal]. None of these works consider the special requirements of cryptographic schemes.
@@
Accelerating polynomial multiplication within specific algebraic structures is essential for enhancing the performance of various cryptographic schemes. In this work, we explore the use of Tensor Processing Units (TPUs), to accelerate polynomial multiplication and, consequently, cryptographic operations. We address two primary challenges in this endeavor: (i) employing the Residue Number System (RNS) to manage large coefficients unsupported by TPUs, and (ii) utilizing a divide-and-conquer approach to handle high-degree polynomials that exceed hardware capacity.
For future work, we intend to focus on the following three key areas: (i) Explore the application of Tensor Processing Units (TPUs) for NTT-based polynomial multiplication; (ii) Enhance the current prototype by optimizing the hardware configuration tailored to specific polynomial rings; and (iii) Design and validate end-to-end cryptographic schemes utilizing TPUs, while investigating potential optimization strategies.